# GOALS-JWST: The Warm Molecular Outflows of the Merging Starburst Galaxy NGC 3256

### v1

@@ 4.2.2 Secondary H Gas Component
As shown in Figure 3 and the second and fourth rows of Figure 4, we detect secondary H components that are collimated and symmetric around the S nucleus in the N-S direction. Since these secondary components are spectrally resolved, we define their velocity as the offset of the peak of the component to the peak of the main emission line component. Based on our F-test detection threshold, they are detected out to a projected distance of 650 and 500 pc to the north and south, beyond which their emission is too low to satisfy the F-test. The northward, receding cone grows wider with increasing distance from the S nucleus (closer to the N nucleus), although it appears offset from the S nucleus and there is a noticeable shift eastwards. The spatial structure of southwards, approaching gas appears truncated, particularly in the S(1) emission, however this is likely due to the secondary component falling below our detection threshold (see Section 3). In both the approaching and receding gas, as seen in the velocity maps of Figure 4, the structure of the secondary component is generally co-spatial with the cold CO gas seen in ALMA observations (Sakamoto et al., 2014). At shorter wavelengths (H S(7) 5.5 m), the secondary component is detected closer to the southern nucleus, to about 80 pc, which can be seen in the second row of Figure 4.
@@
The kinematics of the secondary component show maximum projected velocities of +350 km s in the redshifted gas and 290 km s in the blueshifted gas. The velocities appear relatively constant out to the edges of the detected emission, where the emission falls below our detection threshold about 600 pc away from the nucleus. The velocity dispersion is greatest, up to 140 km s, 400 pc south of the S nucleus, and is generally constant ( 90 km s) elsewhere.
4.3 Temperatures and Masses of the Secondary H Gas Component
@@
The extensive wavelength coverage of JWST, combined with its unprecedented sensitivity, allows us to obtain robust gas masses and temperatures of the warm H gas directly from the rotational H lines between 5–17 m. Here, we calculate the mass of the secondary H gas component using fluxes and power-law temperature modeling of the different transitions of H (Togi & Smith, 2016). H S(8)–S(1) fluxes were obtained using an aperture of radius centered on the peak of each of the blueshifted and redshifted secondary components (see dashed circles in Figure 4). This aperture size was chosen to contain as much of the emission as possible, while staying within the FoV of all MIRI channels. Fluxes were aperture and extinction corrected using the Continuum and Feature Extraction (CAFE222https://github.com/GOALS-survey/CAFE, Marshall et al., 2007, Díaz-Santos et al.; in prep.) software which uses the 9.7 m silicate absorption to estimate the level of extinction. Extinction corrections were generally less than 10 and were slightly higher in the blueshifted gas. Only the H S(3) line at 9.67 m had relatively large corrections, up to 25.
@@
To obtain mass and temperature estimates, we use the power-law model described in Togi & Smith (2016). The power-law index, lower, and upper gas temperatures are the primary model parameters. Since varying the temperature when it is greater than 5000 K has a negligible effect on the quality of the fit and the gas mass, we fixed the upper temperature to 5000 K while the other two parameters were treated as free variables. H line fluxes were used to estimate column densities, whose ratios were then used to create an excitation diagram (see Figure 5). A best-fit model was then fit to the column densities, from which the lower temperature and power-law index were used to calculate the total column density and gas mass.
@@
Fits to the secondary H gas component result in a gas temperature of 26319 K and 27512 K, and a power law index of 4.540.11 and 5.890.13 for the blueshifted and redshifted gas, respectively. This results in a warm (250 K) molecular gas mass (including a heavy element correction factor of 1.36) of = (3.50.7)10 and = (5.40.9)10. Since these apertures do not fully encompass the full extent of the secondary component, the masses calculated here are likely lower limits of the total mass.
@@
To account for the gas mass missed by our apertures, we assumed the power-law indices as calculated from the extractions are constant throughout the secondary component gas. Since channel 3 has the largest FoV in our analysis and is thus likely to best trace the full extent of the secondary component, we opted to use the two H lines, S(1) and S(2), that are present in this channel. For both S(1) and S(2), we found that the apertures encompass about 60 and 65 of the total detected flux in the blue and redshifted gas, respectively. This implies that the molecular mass could be as high as = 5.810 and = 8.310. However, these calculations only include the S(8)–S(1) transitions, and not the H S(0) 28.22 m line which falls outside the wavelength coverage of our MIRI/MRS data. The inclusion of this line tends to lower the estimated temperature which, in turn, would result in an increase of the calculated gas mass. Because of this, the masses listed above should be considered lower limits.
@@
To compare the S(1) and S(7) maps on the same angular resolution, we first convolved the S(7) flux map (see second row, first panel of Figure 4). The convolution was performed using a Gaussian kernel, which reduced the resolution of the S(7) map (FWHM=) to that of the S(1) map (FHWM=). To account for the difference in pixel size of the two maps, we subsequently rebinned the convolved S(7) map to the larger pixel size of the S(1) map. We then calculated the S(7)/S(1) ratios by dividing the two maps, where pixels with reliable flux measurements in both maps were kept.
@@
The resulting ratio map is shown in Figure 6, where higher line ratios indicate a larger warm gas mass fraction. We find the largest warm gas fractions close to the S nucleus, with the ratios decreasing at larger radii. Our calculated power law indices for the blueshifted and redshifted outflows, 4.54 and 5.89, suggest a higher warm gas fraction in the blueshifted outflow, which is also evident from the temperature distribution. In the redshifted, northwards gas, we measure ratios of 0.6 at a distance of 300 pc from the S nucleus. This decreases to 0.3 at 600 pc, where the H fluxes reach their peak in the redshifted gas. For a given distance from the S nucleus, we find slightly higher average ratios in the blueshifted gas, 1.3 at 300 pc, compared to the redshifted gas, 0.6. Unfortunately, we are not able to assess the degree to which the temperature changes beyond 450 pc from the S nucleus, due to the secondary component expanding beyond the MIRI channel 1 FoV.
4.4 Gas Excitation of the ISM
@@
The transitions of H, along with ionized hydrogen recombination lines and [Fe ii], can be used to analyze the gas excitation of the ISM. H can be excited through thermal processes, such as collisional excitation, as well as non-thermal processes, such as UV (6–13.6 eV) photon absorption. In each case, flux ratios of H lines can be used to help distinguish the different mechanisms (Guillard et al., 2009; Mazzalay et al., 2013; Petric et al., 2018; U et al., 2019). Iron can also serve as a tracer of shocks, since it is mostly embedded in dust grains and released only when these grains are destroyed through thermal sputtering caused by shocks (Storchi-Bergmann et al., 2009; Riffel et al., 2013; Inami et al., 2013). Due to the relative abundance of these lines, trends between H/H and [Fe ii]/H have been used to help distinguish excitation sources (Larkin et al., 1998; Riffel et al., 2013; Colina et al., 2015; U, 2022; Bianchin et al., 2023).
@@
In the left panel of Figure 7, we explore the spatial variations of gas excitation by investigating the relation between the ratios [Fe ii] 1.25 m/Pa versus H 1-0 S(1) 2.12 m/Br in the nuclear region of the southern galaxy. Here, line fluxes were calculated from the sum of both the main and secondary components, if present. To show the spatial distribution of the line ratios, we color-code every spaxel in accordance to a linear fit to the ratios, producing an excitation map covering the full NIRSpec FoV, as shown in the right panel of Figure 7. The excitation structure is complex, but there is a noticeable peak in the line ratios to the south of the S nucleus. To the north of the nucleus, there is a broad, high excitation structure that extends westwards. The lowest line ratios are located in the S nucleus and in regions coincident with multiple clumps that are detected in NIRCam imaging (see Linden et al., in prep.). The implications of these regions and corresponding line ratios are discussed in Section 5.3.
@@ 5 Discussion
The high velocities and structure of the secondary H emission line component are most consistent with outflowing gas, and we hereafter refer to the secondary component as the outflow component. Although outflowing warm H2 gas has previously been identified (e.g., Emonts et al., 2017; U et al., 2019), this is the first time that the warm H gas in the MIR range is spatially and spectrally resolved.
It is noteworthy, however, that we do not detect a secondary H component in or immediately around the N nucleus. Specifically, our MIR spectra do not show H lines with sufficient asymmetry to pass our detection thresholds (see Section 3) in the regions where the cold molecular outflow from the N nucleus is detected (Sakamoto et al., 2014; Michiyama et al., 2018). As such, for the remainder of this article, we will only reference and analyze the outflows associated with the S nucleus.
@@ 5.1 Outflow Characteristics
The outflows are detected out to a distance of 650 and 500 pc (deprojected: 690 and 530 pc) to the north and south of the nucleus. The regions with the highest outflow emission are spatially coincident with strong emission in the main H component and cold CO gas. Moreover, distribution of the HCN and HCO gas, which trace the denser gas, show enhanced emission in these outflow regions (Michiyama et al., 2018). As such, these H flux peaks could be caused by the outflowing gas impacting a dense region in the ISM.
The kinematics of the outflows show maximum projected velocities to be +350 km s to the north and 290 km s to the south. To obtain intrinsic velocities, we adopt the outflow inclination as calculated by Sakamoto et al. (2014) and Emonts et al. (2014), . Deprojecting the outflows, v/sin, we obtain maximum intrinsic velocities of +1020 and 850 km s, respectively (where v= 750 km s, averaged across all regions of the outflows).
@@ 5.2 Outflow Mass Comparisons of MIR to NIR and Submillimeter Measurements
In Section 4.3, we calculated the mass of the warm outflowing gas to be = (3.50.7)10 and = (5.40.9)10 for the blueshifted and redshifted outflows in the emission peak regions. We can compare these masses to that of the hotter (2000 K) gas observed at NIR wavelengths. To measure this hotter component, we use the H 1-0 S(1) 2.12 m outflow flux map created by Emonts et al. (2014), which was derived from K-band VLT/SINFONI observations (program ID: 078.B-0066(A), PI: Luis Colina). Measuring the flux within the same apertures used to calculate the warm MIR gas mass and following the methods described in Mazzalay et al. (2013) and Emonts et al. (2014), we obtain mass measurements of 33060 and 20035 for the blue and redshifted outflows, respectively. This combined gas mass of 530 is almost 210 times lower than the warm MIR component, (8.91.1)10.
@@
In order to assess the contribution of the warm gas to the total multiphase component, we can use molecular CO to estimate the mass of the outflowing cold H gas. To do this, we opted to use archival CO (2-1) ALMA data (project code: 2015.1.00902.S, PI: Yiping Ao), which has higher spatial resolution than the CO (1-0)333The spatial resolution of the CO (2-1) data is while the CO (1-0) data is . map of Sakamoto et al. (2014). Using the same two r= apertures used in our H mass calculations, we extracted spectra of the CO (2-1) emission. Based on a two-component Gaussian fit, we measured line fluxes of 20.62.1 and 8.61.9 Jy km s with velocity offsets of 150 km s and 250 km s for the blueshifted and redshifted gas, respectively. These are consistent with the mean velocities measured in the warm H outflow, 170 km s and 240 km s. Following Michiyama et al. (2018), we set L/L as unity. For the CO to H conversion factor, =, we assume a value of 0.8 (K km s pc), which is typically used for ULIRGs (Downes & Solomon, 1998; Bryant & Scoville, 1999). This value is 5.4 times lower than the galactic value of 4.32. This is because ULIRGs, particularly those involved in mergers, typically have increased CO linewidths (and thus greater L) caused by the heightened turbulence and temperatures induced by the outflows and mergers. From these assumptions, we obtained = (1.60.2)10 and = (6.81.5)10, which implies the combined outflowing mass is = (2.30.3)10.
@@
While we assumed a value of 0.8 for , Cicone et al. (2018) calculated to be 2.1 where molecular outflows are present. This in turn would increase our cold gas mass estimate to = (6.00.8)10. More recently, Pereira-Santaella et al. (2024) estimated of various regions around the outflows in NGC 3256 utilizing the CO rovibrational band at 4.67 m. From their two-component model, they report to be 0.62. Setting as 0.62 reduces our combined cold gas mass to (1.80.3)10, still consistent with the mass calculated using the canonical conversion factor for local ULIRGs. Hereafter, we adopt 2.310 for the reminder of our analysis.
@@
Combining the redshifted and blueshifted outflows, the / gas mass fraction is about 4. As mentioned previously, we note that our data set does not include the S(0) 28.22 m line which tends to lower the estimated temperature, thus increasing the / fraction. Higdon et al. (2006) measured typical warm gas mass fractions in observed ULIRGs to be slightly less than 1 (as calculated from the S(1) 17.0 m to S(3) 9.67 m line fluxes). However, they found that the inclusion of S(0) could increase / to tens of percent. While the warm mass fraction may depend heavily on the inclusion of S(0), our estimates for the warm mass fraction in the outflows are comparable to those measured in other local active galaxies (Rigopoulou et al., 2002).
@@
In Section 5.1, we calculated the maximum intrinsic outflow velocity to be 10 km s, and maximum deprojected distance to be 700 pc. Following this, the outflow time scale, , is estimated to be yr. Assuming a conical geometry, this results in a time-averaged outflow mass rate of = 1.3 M yr. If a significant portion of the outflow mass is in the cold component, then the outflow mass rate could be as high as = 33 M yr, which is slightly lower than the 48 M yr reported by Sakamoto et al. (2014). Regardless, these values are comparable to rates seen in other Seyferts and ULIRGs (García-Burillo et al., 2014; Cicone et al., 2014). These rates, particularly the cold gas mass rate, are high compared to estimates of the star formation rate (SFR). In regions around the S nucleus, Emonts et al. (2014) and Michiyama et al. (2018) estimated the SFR to be 1 - 6.8 M yr based on hydrogen recombination line calibrations. For the cold molecular outflow, this leads to a mass loading factor, /SFR, of 5 or greater. Amongst starburst galaxies, mass loading factors are typically around unity while AGN-dominated sources tend to be higher () (Cicone et al., 2014).

### v3

@@ 4.2.2 Secondary H2 Gas Component
As shown in Figure 3 and the second and fourth rows of Figure 4, we detect secondary H2 components that are collimated and symmetric around the S nucleus in the north-south direction. These secondary components are spectrally resolved, and we define their velocity as the offset from the systemic velocity of the S nucleus, as calculated from the 2.3 m CO absorption bands. Based on our F-test detection threshold, they are detected out to a projected distance of 650 and 500 pc to the north and south, respectively, beyond which their emission is too low to satisfy the F-test. The northward, receding cone grows wider with increasing distance from the S nucleus (closer to the N nucleus), although it appears offset from the S nucleus and there is a noticeable shift eastwards. The spatial structure of southwards, approaching gas appears truncated, particularly in the S(1) emission, however this is likely due to the secondary component falling below our detection threshold (see Section 3). In both the approaching and receding gas, as seen in the velocity maps of Figure 4, the structure of the secondary component is generally cospatial with the cold CO gas seen in ALMA observations (Sakamoto et al., 2014). At shorter wavelengths (H2 S(7) 5.5 m), the secondary component is detected closer to the S nucleus, to about 80 pc, which can be seen in the second row of Figure 4.
@@
The kinematics of the secondary component show maximum projected velocities of +340 km s-1 in the redshifted gas and 310 km s-1 in the blueshifted gas. The velocities appear relatively constant out to the edges of the detected emission, where the emission falls below our detection threshold about 600 pc away from the nucleus. The velocity dispersion is greatest, up to 140 km s-1, 400 pc south of the S nucleus, and is generally constant ( 90 km s-1) elsewhere.
To compare this to the warmer gas phase, which can be traced by H2 2.12 m emission, we refer to VLT/SINFONI data which have a larger FOV than JWST/NIRSpec. Here, peak projected velocities of +270 km s-1 and 240 km s-1, and a peak velocity dispersion of 190 km s-1 are measured in H2 2.12 m emission (Emonts et al., 2014), both of which are lower than the MIR component. The locations of these peaks, however, are cospatial.
| H2 Line | FluxBlue | FluxRed,N | FluxRed,S |
| (m) | (10-18 W m-2) | ||
| (1) | (2) | (3) | (4) |
| S(8) 5.05 | 1.300.10 | 0.180.02 | — |
| S(7) 5.51 | 6.970.59 | 3.070.24 | 1.700.14 |
| S(6) 6.11 | 2.560.21 | 0.740.06 | — |
| S(5) 6.91 | 14.131.07 | 6.510.49 | 4.190.31 |
| S(4) 8.02 | 4.310.34 | 3.340.25 | 1.670.13 |
| S(3) 9.66 | 8.880.68 | 6.640.50 | — |
| S(2) 12.28 | 3.750.31 | 4.680.41 | 2.120.19 |
| S(1) 17.03 | 5.720.46 | 8.460.74 | 3.070.27 |
| Tlow (K) | 26319 | 27512 | 31322 |
| n | 4.540.11 | 5.890.13 | 5.320.22 |
| Mout (10) | 3.50.7 | 5.40.9 | 1.50.5 |
Note. — Columns: (1) H2 emission lines with rest-frame wavelengths. (2) Flux within an aperture of radius of the blueshifted secondary component, centered on its peak emission. The listed errors are representative of uncertainties in the fitting process. (3–4) Same as column 2, but for the redshifted secondary component within the northern and southern regions.
—Bottom rows: The final three rows list the lower gas temperature, power-law index, and mass of the secondary component gas for each region.
4.3 Temperatures and Masses of the Secondary H2 Gas Component
@@
The extensive wavelength coverage of JWST, combined with its unprecedented sensitivity, allows us to obtain robust gas masses and temperatures of the warm H2 gas directly from the rotational H2 lines between 5–17 m. Here, we calculate the mass of the secondary H2 gas component using fluxes and power-law temperature modeling of the different transitions of H2 (Togi & Smith, 2016). Aperture-corrected H2 S(8)–S(1) fluxes were obtained using the r = apertures centered on the peak of the blueshifted and redshifted secondary components (see the blue and red dashed circles in Figure 4 and Table 1). This aperture size was chosen to contain as much of the emission as possible, while staying within the FOV of all MIRI channels. The third aperture was used to sample the redshifted gas closer to the S nucleus (see orange dotted circle in Figure 4). Its position is at a similar distance to the S nucleus as the aperture sampling the blueshifted gas so as to allow for a direct comparison of outflowing gas conditions. However, secondary, redshifted H2 S(3), S(6) and S(8) emission were not included in the fit due to not passing our S/N 5 detection threshold. Extinction correction was done using the 9.7 m silicate absorption, where corrections were generally less than 10 and were slightly higher in the blueshifted gas. Only the H2 S(3) line at 9.66 m had relatively large corrections, up to 25.
@@
To obtain mass and temperature estimates, we use the power-law model described in Togi & Smith (2016). The power-law index, lower, and upper gas temperatures are the primary model parameters. Since varying the temperature when it is greater than 5000 K has a negligible effect on the quality of the fit and the gas mass, we fixed the upper temperature to 5000 K while the other two parameters were treated as free variables. The model assumes local thermodynamic equilibrium and samples a range of power-law indices and lower gas temperatures, the latter of which determines the ortho-to-para ratio based on Figure 1 of Sternberg & Neufeld (1999). H2 line fluxes were used to estimate column densities, whose ratios were then used to create an excitation diagram (see Figure 5). A best-fit model was then fit to the column densities, from which the lower temperature and power-law index were used to calculate the total column density and gas mass.
@@
Fits to the secondary H2 gas component result in a lower gas temperature of 26319 K, 27512 K, and 31322 K (ortho-to-para ratio of 3), and a power-law index of 4.540.11, 5.890.13, and 5.320.22 for the blueshifted, redshifted (north), and redshifted (south) gas, respectively. In other local LIRGs, typical power law indices range from 3.4 to 5.4 based on Spitzer/IRS data (Togi & Smith, 2016), which are consistent with our values. Our fits result in a warm molecular gas mass (including a heavy element correction factor of 1.36) of = (3.50.7)10, = (5.40.9)10, and = (1.50.5)10. Since these apertures do not fully encompass the full extent of the secondary component, the masses calculated here do not represent the total gas mass associated with the secondary component.
@@
To account for the gas mass missed by our apertures, we assumed the power-law indices as calculated from the extractions are constant throughout the secondary component gas. Since channel 3 has the largest FOV in our analysis and is thus likely to best trace the full extent of the secondary component, we opted to use the two H2 lines, S(1) and S(2), that are present in this channel. For both S(1) and S(2), we found that the apertures at the emission peak encompasses about 60 of the total detected flux in the blueshifted gas. Assuming a constant power-law index of 4.54 and a temperature of 263 K, this implies that the molecular mass could be as high as = (5.81.2)10. For the redshifted gas, we found that the northern and southern apertures encompass roughly 65 and 20 of the total detected redshifted flux in both S(1) and S(2). Extrapolating using the northern region values, n=5.89 and Tlow=275 K, leads to a total estimated gas mass of = (8.31.4)10. Similarly, extrapolating using the southern region values, n=5.32 and Tlow=313 K, results in a total gas mass of = (7.92.6)10. These two estimates are in good agreement, and we opt to use (8.31.4)10 since the aperture used encompasses a larger fraction of the total emission.
However, the above calculations only include the S(8)–S(1) transitions, and not the H2 S(0) 28.22 m line, which falls outside the wavelength coverage of our MIRI/MRS data. The inclusion of this line tends to lower the estimated temperature, which, in turn, would result in an increase of the calculated gas mass. While the masses listed above should be considered lower limits, we use = (5.81.2)10 + (8.31.4)10 = (1.40.2)10 as the total gas mass of the secondary component for our analysis.
@@
To compare the S(1) and S(7) maps on the same angular resolution, we first convolved the S(7) flux map (see second row, first panel of Figure 4). The convolution was performed using a Gaussian kernel, which reduced the resolution of the S(7) map (FWHM=) to that of the S(1) map (FHWM=). To account for the difference in pixel size of the two maps, we subsequently rebinned the convolved S(7) map to the larger pixel size of the S(1) map. We then calculated the S(7)/S(1) ratios by dividing the two maps, where pixels with S/N 5 in both maps were kept (see Section 3).
@@
The resulting ratio map is shown in Figure 6, where higher line ratios indicate a larger warm gas mass fraction. We find the largest warm gas fractions close to the S nucleus, with the ratios decreasing at larger radii. Our calculated power-law indices for the blueshifted (4.54) and redshifted (5.32 and 5.89) outflows suggest a higher warm gas fraction in the blueshifted outflow, which is also evident from the temperature distribution. In the redshifted, northwards gas, we measure S(7)/S(1) ratios of 0.6 at a distance of 300 pc from the S nucleus. This decreases to 0.3 at 600 pc, where the H2 fluxes reach their peak in the redshifted gas. For a given distance from the S nucleus, we find slightly higher average ratios in the blueshifted gas, 1.3 at 300 pc, compared to the redshifted gas, 0.6. Unfortunately, we are not able to assess the degree to which the temperature changes beyond 450 pc from the S nucleus, due to the secondary component expanding beyond the MIRI channel 1 FOV.
4.4 Gas Excitation of the Interstellar Medium
@@
The transitions of H2, along with ionized hydrogen recombination lines and [Fe II], can be used to analyze the gas excitation of the ISM. H2 can be excited through thermal processes, such as collisional excitation, as well as nonthermal processes, such as UV photon absorption. In each case, flux ratios of H2 lines can be used to help distinguish the different mechanisms (Guillard et al., 2009; Mazzalay et al., 2013; Petric et al., 2018; U et al., 2019). Iron can also serve as a tracer of shocks, since it is mostly embedded in dust grains and released only when these grains are destroyed through thermal sputtering caused by shocks (Storchi-Bergmann et al., 2009; Riffel et al., 2013; Inami et al., 2013). Due to the relative abundance of these lines, trends between H2/H and [Fe II]/H have been used to help distinguish excitation sources (Larkin et al., 1998; Riffel et al., 2013; Colina et al., 2015; U, 2022; Bianchin et al., 2024).
@@
In the left panel of Figure 7, we explore the spatial variations of gas excitation by investigating the relation between the ratios [Fe II] 1.25 m/Pa versus H2 1-0 S(1) 2.12 m/Br in the nuclear region of the southern galaxy. The purpose of this relation is to identify enhanced emission which could be caused by shocks, and we thus use the total emission of each line. To show the spatial distribution of the line ratios, we color code every spaxel in accordance to a linear fit to the ratios, producing an excitation map covering the full NIRSpec FOV, as shown in the right panel of Figure 7. The excitation structure is complex, but there is a noticeable peak in the line ratios to the south of the S nucleus. The median values of the ratios in this region are [Fe II]/Pa = -0.06 and H2/Br = 0.89, compared to -0.17 and 0.16 for regions outside the outflow (see blue and gray stars in left panel of Figure 7). To the north of the nucleus, there is a broad structure with slightly higher excitation, median ratios of [Fe II]/Pa = -0.16 and H2/Br = 0.64 (red star in Figure 7 left), that extends westwards. The lowest line ratios are located in the S nucleus and in regions coincident with multiple clumps that are detected in NIRCam imaging (Linden et al., 2024). The implications of these regions and corresponding line ratios are discussed in Section 5.3.
@@ 5 Discussion
The high velocities and structure of the secondary H2 emission-line component are most consistent with outflowing gas, and we hereafter refer to the secondary component as the outflow component. Although outflowing warm H2 gas has previously been identified (e.g., Emonts et al., 2017; U et al., 2019), this is the first time that the warm, outflowing H2 gas in the MIR range is spatially and spectrally resolved.
It is noteworthy, however, that we do not detect a secondary H2 component in or immediately around the N nucleus. Specifically, our MIR spectra do not show H2 lines with sufficient asymmetry to pass our detection thresholds (see Section 3) in the regions where the cold molecular outflow from the N nucleus is detected (Sakamoto et al., 2014; Michiyama et al., 2018). As such, for the remainder of this article, we will only reference and analyze the outflows associated with the S nucleus.
@@ 5.1 Outflow Characteristics
The outflows are detected out to a distance of 650 and 500 pc to the north and south of the nucleus. To obtain deprojected distances, we adopt the outflow inclination as calculated by Sakamoto et al. (2014) and Emonts et al. (2014), , which results in intrinsic distances of 690 and 530 pc. The regions with the highest outflow emission are spatially coincident with strong emission in the main H2 component and cold CO gas. Moreover, the distribution of the HCN and HCO+ gas, which trace the denser gas, show enhanced emission in these outflow regions (Michiyama et al., 2018). As such, these H2 flux peaks could be caused by the outflowing gas impacting a dense region in the ISM.
The kinematics of the outflows show maximum projected velocities to be +340 km s-1 to the north and 310 km s-1 to the south. Deprojecting the outflows, vout/sin, we obtain maximum intrinsic velocities of +990 and 910 km s-1, respectively (where vout,avg= 730 km s-1, averaged across all regions of the outflows).
@@ 5.2 Outflow Mass Comparisons of MIR to NIR and Submillimeter Measurements
In Section 4.3, we calculated the total mass of the warm outflowing gas to be = (5.81.2)10 and = (8.31.4)10 for the blueshifted and redshifted outflows, respectively. We can compare these masses to that of the hotter (2000 K) gas observed at NIR wavelengths. To measure this hotter component, we use the H2 1-0 S(1) 2.12 m outflow flux map created by Emonts et al. (2014), which was derived from K-band VLT/SINFONI observations (program ID: 078.B-0066(A), PI: Luis Colina). Measuring the flux within the same apertures used to calculate the warm MIR gas mass and following the methods described in Mazzalay et al. (2013) and Emonts et al. (2014), we obtain mass measurements of 33060 and 20035 for the blue and redshifted outflows, respectively. This combined gas mass of 530 is over 2.5103 times lower than the warm MIR component, (1.40.3)10.
@@
In order to assess the contribution of the warm gas to the total multiphase component, we can use molecular CO to estimate the mass of the outflowing cold H2 gas. To do this, we opted to use archival CO (2-1) ALMA data (project code: 2015.1.00902.S, PI: Yiping Ao), which has higher spatial resolution than the CO (1-0) map222The spatial resolution of the CO (2-1) data is while the CO (1-0) data is . of Sakamoto et al. (2014). Using the same two r= apertures for our H2 mass calculations, we extracted spectra of the CO (2-1) emission. Based on a two-component Gaussian fit, we measured line fluxes of 20.62.1 and 8.61.9 Jy km s-1 with velocity offsets of 150 km s-1 and 250 km s-1 for the blueshifted and redshifted gas, respectively. These are consistent with the mean velocities measured in the warm H2 outflow, 170 km s-1 and 240 km s-1. Following Michiyama et al. (2018), we set LCO(2-1)/LCO(1-0) as unity. For the CO to H2 conversion factor, =, we assume a value of 0.8 (K km s-1 pc2)-1, which is typically used for ULIRGs (Downes & Solomon, 1998; Bryant & Scoville, 1999). This value is 5.4 times lower than the galactic value of 4.32. This is because ULIRGs, particularly those involved in mergers, typically have increased CO line widths (and thus greater LCO) caused by the heightened turbulence and temperatures induced by the outflows and mergers (Togi & Smith, 2016). From these assumptions, we obtained = (1.60.2)10 and = (6.81.5)10, which implies the combined outflowing mass is = (2.30.2)10.
@@
While we assumed a value of 0.8 for , Cicone et al. (2018) calculated to be 2.1 where molecular outflows are present. This in turn would increase our cold gas mass estimate to = (6.00.8)10. More recently, Pereira-Santaella et al. (2024) estimated of various regions around the outflows in NGC 3256 utilizing the CO rovibrational band at 4.67 m. From their two-component model, they report to be 0.62. Setting as 0.62 reduces our combined cold gas mass to (1.80.3)10, still consistent with the mass calculated using the canonical conversion factor for local ULIRGs. Hereafter, we adopt 2.310 for the remainder of our analysis.
@@
Combining the redshifted and blueshifted outflows, the / gas mass fraction is about 6. As mentioned previously, we note that our data set does not include the S(0) 28.22 m line which tends to lower the estimated temperature, thus increasing the / fraction. Higdon et al. (2006) measured typical warm gas mass fractions in observed ULIRGs to be slightly less than 1 (as calculated from the S(1) 17.0 m to S(3) 9.66 m line fluxes). However, they found that the inclusion of S(0) could increase / to tens of percent. While the warm mass fraction may depend heavily on the inclusion of S(0), our estimates for the warm mass fraction in the outflows are comparable to those measured in other local active galaxies, which generally range from 1–10, though it can reach up to 35 in some cases (Rigopoulou et al., 2002).
@@
In Section 5.1, we calculated the maximum intrinsic outflow velocity to be 103 km s-1, and maximum deprojected distance to be 700 pc. Following this, the outflow time scale, , is estimated to be yr. Assuming a conical geometry, this results in a time-averaged outflow mass rate of = 2.00.8 M⊙ yr-1. If a significant portion of the outflow mass is in the cold component, then the outflow mass rate could be as high as = 33 M⊙ yr-1, which is slightly lower than the 48 M⊙ yr-1 reported by Sakamoto et al. (2014), though our value here should be considered a lower limit due to our use of a r = aperture. Regardless, these values are comparable to rates seen in other Seyferts and ULIRGs (García-Burillo et al., 2014; Cicone et al., 2014). These rates, particularly the cold gas mass rate, are high compared to estimates of the SFR. In regions around the S nucleus, Emonts et al. (2014) and Michiyama et al. (2018) estimated the SFR to be 1 - 6.8 M⊙ yr-1 based on hydrogen recombination line calibrations. For the cold molecular outflow, this leads to a mass loading factor, /SFR, of five or greater. Amongst starburst galaxies, mass loading factors are typically around unity while AGN-dominated sources tend to be higher () (Cicone et al., 2014).